OpenStudy (anonymous):

Find (d^2 * y)/(dx^2) for y = e^x^2 (No typos, just a really bizarre one.)

OpenStudy (abb0t):

\(\frac{d^2y}{dx^2}\) is merely using chain rule and then product rule..

OpenStudy (abb0t):

\(y' = e^{x^2} \times(2x)\) and then \(y'' = (2x)(e^{x^2})\)

OpenStudy (abb0t):

solve for y'', i just left it as is, i didn't solve it.

OpenStudy (anonymous):